OpenStudy (anonymous):

What is the best estimate for the answer to 38.94 + 65.43?

OpenStudy (anonymous):

PLEASE I NEED HELP

OpenStudy (gabbyalicorn):

38.94+65.43 =38.94+65.43 =104.37 Well, the real answer is 104.37, so maybe you could round it to just 104. :) Another way to do it is rounding 38.94 to 40 & 65.43 to 65. 40 + 64=104. :)

OpenStudy (gabbyalicorn):

Do you understand? Medal if helped?

OpenStudy (anonymous):

YES ANYTHING